OpenStudy (anonymous):

limit as x approaches 0 x^8-1/x-1

OpenStudy (solomonzelman):

\[\lim_{x \rightarrow 0}\frac{x^8-1}{x-1}\]

OpenStudy (solomonzelman):

you can plug in x=0 right away.

OpenStudy (solomonzelman):

because when you plug in x=0, you don't get any undefined values.

OpenStudy (solomonzelman):

\[\lim_{x \rightarrow 0}\frac{x^8-1}{x-1}~=~\frac{(0)^8-1}{(0)-1}=\frac{-1}{-1}=1\]

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/q1p7e3ybo3

OpenStudy (solomonzelman):

I put in 0.05 and not 0, so that it would be actually not touching the x-axis.

OpenStudy (solomonzelman):

in other words, when limits of x approach 0, the function approaches 1. That is a point (0,1) And this is where it looses continuity.

OpenStudy (solomonzelman):

if you need more info.... tell me. and slope is always increasing, but there are 2 concavity changes

OpenStudy (zarkon):

are you sure you wanted \(x\to0\) ? \(x\to1\) is a more interesting problem

OpenStudy (isaiah.feynman):

Hmmm @Zarkon you are right!

OpenStudy (solomonzelman):

yes... then you would use L'H"S .. or that would be the easiest approach.

OpenStudy (solomonzelman):

L'Hopital's rule (that's the abbreviation)

OpenStudy (solomonzelman):

\[\lim_{{x~} \rightarrow {~1}}~\frac{x^8-1}{x-1}=\large \lim_{{x~} \rightarrow {~1}}~\frac{\frac{d}{dx}(x^8-1)}{\frac{d}{dx}(x-1)}=\large \lim_{{x~} \rightarrow {~1}}~\frac{8x^7}{1}=\lim_{{x~} \rightarrow {~1}}~8x^7\]

OpenStudy (solomonzelman):

and when you plug in 1 for x, you get 1.

OpenStudy (isaiah.feynman):

8 actually, :P

OpenStudy (solomonzelman):

ohh, man you are right.

OpenStudy (solomonzelman):

LOL, how in the world did I get 1? ty for catching the mistake:)

OpenStudy (isaiah.feynman):

Man, I had make silly mistakes too! I forgot intercepts one time lol.

OpenStudy (isaiah.feynman):

have*

OpenStudy (zarkon):

\[x^8-1=(x^4-1)(x^4+1)=(x^2-1)(x^2+1)(x^4+1)\] \[=(x-1)(x+1)(x^2+1)(x^4+1)\]

OpenStudy (solomonzelman):

yeah.. in my definite integrals I kept getting the wrong answer, not thinking that 1/2 - 1/4 is 1/4, not 1/2 :)

OpenStudy (solomonzelman):

I like that Zarkon:)

OpenStudy (isaiah.feynman):

Yes yes!! Zarkon just what I WAS PONDERING

OpenStudy (solomonzelman):

he s a better latex typer :P

OpenStudy (zarkon):

\[\lim_{x\to1}\frac{x^8-1}{x-1}=\lim_{x\to1}\frac{(x-1)(x+1)(x^2+1)(x^4+1)}{x-1}\] \[=\lim_{x\to1}(x+1)(x^2+1)(x^4+1)=(1+1)(1^2+1)(1^4+1)=2\times2\times2=8\]